OpenStudy (anonymous):

A manufacturer produces bolts of a fabric with a fixed width. A quantity q of this fabric (measured in yards) that is sold is a function of the selling price p (in dollars per yard), so we can write q=f(p). Then, the total revenue earned with selling price p is R(p)=pf(p). Find R′(20), given f(20)=10000, and f′(20)=−350. (What does this mean?)

OpenStudy (anonymous):

R'(p)=p*f(p) \[R'(p) = \frac{d}{dp}R(p)\] It's asking you to find the derivative of R(p) at p=20 You need to apply the product rule.